OpenStudy (anonymous):

tan^x-sec^4x = 1 - 2sec^2x

OpenStudy (anonymous):

...is this supposed to be \[\tan(x) - \sec^4(x) = 1-2\sec^2(x)\]?

OpenStudy (anonymous):

oh no. oops. its tan^4(x) - sec^4(x) = 1 - 2 sec^2(x)

myininaya (myininaya):

ok you have the difference of squares here a^2+b^2=(a-b)(a+b)

OpenStudy (anonymous):

okay so that would make it (tan^2x-sec^2x) (tan^2x+sec^2x)

myininaya (myininaya):

so we have (tan^2x)^2-(sec^2x)^2=[(tanx)^2+(secx)^2][(tanx)^2-(secx)^2]

myininaya (myininaya):

yes and 1=(secx)^2-(tanx)^2

myininaya (myininaya):

so we have -((tanx)^2+(sec^2))((secx)^2-(tanx)^2)

myininaya (myininaya):

-((tanx)^2+(secx)^2)(1)

myininaya (myininaya):

ok the other side is in terms of sec so put this side in terms of sec

OpenStudy (anonymous):

ok thanks

myininaya (myininaya):

you got it?

myininaya (myininaya):

this problem is basically done

myininaya (myininaya):

so we have -((secx)^2-1+(secx)^2)=-(2(secx)^2-1)=1-2(secx)^2

myininaya (myininaya):

which is the other side of the equation